# Discovering gravitational waveform distortions from lensing: a deep dive into GW231123

### v1

@@ Abstract
Gravitational waves (GWs) are unique messengers as they travel through the Universe without alteration except for gravitational lensing. Their long wavelengths make them susceptible to diffraction by cosmic structures, providing an unprecedented opportunity to map dark matter substructures. Identifying lensed events requires the analysis of thousands to millions of simulated events to reach high statistical significances. This is computationally prohibitive with standard GW parameter estimation methods. We build on top of state-of-the-art neural posterior algorithms to accelerate the lensed inference from CPU days to minutes with DINGO-lensing. We showcase its capabilities by reanalyzing GW231123, the most promising lensed candidate so far, and find that its statistical significance cannot exceed $4\sigma$. We observe that 8% of GW231123-like nonlensed simulations favor lensing, which could be explained by the self-similarity of short-duration signals. Still, 58% of GW231123-like lensed simulations have larger support for lensing, showing that higher detection statistics are possible. Although GW231123 exposes the challenges of claiming the first GW lensing detection, our deep-learning methods have demonstrated to be powerful enough to enable the upcoming discovery of lensed GWs.
@@
GW parameter estimation has recently experienced a revolution, thanks to deep-learning techniques. In particular, neural posterior estimation [40, 41, 42] is able to drastically reduce the inference time from order of days down to minutes by avoiding the expensive likelihood evaluations needed by conventional methods [43]. This capability enables the assessment of statistical significance for subtle astrophysical effects, such as orbital eccentricity [44], which was previously not feasible due to the prohibitive computational cost of traditional sampling algorithms. In this Letter, we demonstrate that deep-learning techniques are essential for the discovery of lensed GWs. Building on top of state-of-the-art algorithm, DINGO [40, 42, 45, 46, 47], we develop DINGO-lensing [48], which we use to thoroughly reanalyze GW231123 and assess its statistical significance. Our results and methods lay out the detections strategies of lensed GWs for upcoming observing runs.
Gravitational waveform distortions from lensing.—Starting from an emitted GW signal $h(t)$, the observed lensed waveform $h_{\rm L}(t)$ is obtained by propagating $h(t)$ through the curved spacetime defined by the lens. Assuming linear perturbations, weak-field gravity, and a lens much smaller than the source–observer distance, this propagation is conveniently solved in Fourier space. Lensing acts like a transfer function between the emitted and observed signals, $F(f)=\tilde{h}_{\rm L}(f)/\tilde{h}_{\rm U}(f)$, where the amplification function $F(f)$ is defined by the Kirchhoff diffraction integral:
@@
which integrates the time delay $t_{d}$ accumulated by the lensed signal over all paths from source to observer. Here $\vec{\theta}_{\mathrm{S}}$ is the source position, $\vec{\theta}$ the angular position on the lens plane, and $\tau_{D}=\frac{D_{\rm L}D_{\rm S}}{cD_{\rm LS}}$ is a reference time related to the angular diameter distances to the lens, the source, and between the lens and the observer, respectively. The frequency dependence of $F(f)$ produces the characteristic waveform distortions in $h_{\rm L}(t)$.
The nature of these distortions depends on the lens profile and the source-lens configuration. Recent studies have developed efficient numerical methods to solve the diffraction integral for different lens classes [49, 50, 51]. Observationally, the strongest distortions arise when repeated chirps from a strongly lensed GW source interfere, while in the deep wave-optics regime diffraction effects are mild, with $\lim_{f\to 0}F(f)=1$. This interference regime can be modeled accurately using the stationary phase approximation (SPA) [52], yielding
@@
where $\Delta t\$and $\mu_{\rm rel}\$are the time delay and relative magnification between two repeated signals, respectively. We assume the second chirp has a lower magnification and opposite parity, i.e. $\Delta t\$>0>0, $\mu_{\rm rel}\$<1<1, and a $\pi/2$ phase difference  [^2] Our parametrization differs from the phenomenological model in Ref. [75], where the number of repeated chirps and their phase differences are treated as free parameters.. The overall magnification is omitted, as it is degenerate with luminosity distance. This parametrization encompasses point and fold singularities, as well as many near-cusp configurations [26, 52, 54]. Folds and cusps are the most common caustics in generic lenses [55], while point singularities correspond to axisymmetric lenses such as point-mass lenses or singular isothermal spheres. Moreover, this simple parametrization has been shown to efficiently recover simulated signals lensed by point-mass lenses [48]. Negative-parity single images, which may cause small additional distortions [56, 57], are excluded here and left for future work.
Under these assumptions, we define a lensed event as one where the analysis yields $\mu_{\rm rel}\$>0>0 and $\Delta t\$>0>0. More complex cases beyond Eq. (2), such as overlapping images of the same type with a phase different from $\pi/2$ or $\mu_{\rm rel}\$>1>1 due to substructure [13, 14], will be explored in follow-up work.
@@
We train the lensed networks with $\Delta t\subset(0,0.1)$ sec, and $\mu_{\text{rel}}\subset(0,1)$. The nonlensed parameters follow standard choices for the extrinsic parameters and spins (with magnitudes up to 0.99) [48]. Lensed waveforms in DINGO-lensing and bilby are implemented through modwaveforms  [^12] https://github.com/ezquiaga/modwaveforms. We validate the networks as described in [48], testing the recovery of the injected values, comparing the inferences for representative examples with bilby [43], and checking the sampling efficiencies when performing importance sampling [47]. DINGO-lensing  drastically reduces the computing time compared to conventional methods, sampling with neural posterior estimations in seconds and computing evidences with importance sampling in minutes. In comparison, point-mass lens analyses of real events with Gravelamps [49], a code used by LVK, take on average $\sim 50$ CPU days.
Figure 1: Source characterization of GW231123 assuming it is lensed (with DINGO-lensing, in blue) and not lensed (with DINGO, in green), respectively. The posterior distributions using DINGO-lensing and DINGO are compatible with those obtained using bilby. The credible levels quoted here are from the DINGO-lensing analysis.
@@
Figure 2: Inference for the lensing parameters. We compare GW231123 results (filled contours) with nonlensed simulations for similarly heavy binaries (unfilled contours).
Using DINGO-lensing, we train neural networks to learn (non)lensed signals using a numerical relativity surrogate waveform model, NRSur7dq4 [66], with a frequency range of $20,512]~\rm{Hz}$, and an analysis duration of $8~\rm{sec}$  [^13] We set the waveform starting frequency at 0 Hz. This means that the model will generate waveforms as far back in time from the merger as allowed.. NRSur7dq4 is the waveform model with the lowest mismatch compared to some GW231123-like numerical relativity simulations [30]. We train the networks for chirp masses in the range of $30,180]~M_{\odot}$, mass ratios in $[1/6,1]$ and luminosity distances in $0.5,10]~\rm{Gpc}$. The prior distribution of the remaining source parameters follow Table 1 in [48], with extended constraints in the component masses. We use the average noise power spectral density of the LIGO detectors around the time of the event [68], and stop the training at epoch 300 when the loss function plateaus.
@@
Figure 3: Whitened best-fit lensed (solid), unlensed (dash), and nonlensed (dotted) waveforms for GW231123. The inferred time delay between the two chirps $\Delta t$ matches with their instantaneous period near the reference time.
We observe that in many other instances, the recovered $\Delta t$ is multimodal, and the peaks correlate with the period of the waveform. This is due to the self-similarity of short waveforms such as GW231123, for which only a few cycles are observable. We illustrate this using GW231123 in Fig. 3, where we show the whitened best-fit lensed, unlensed (i.e., undoing the lensing amplification from the lensed waveform), and nonlensed (i.e., assuming that there is no lensing) waveforms, respectively  [^15] The whitening uses the estimated noise power spectral density for GW231123 from the LIGO Hanford detector [30].. Our inferred time delay between the two chirps $\Delta t\approx 22$ ms matches with the instantaneous period of the waveforms near the merger at around the reference time $t_{\rm ref}$.
Furthermore, we compare GW231123 Bayes factor with a set of 70,000  GW231123-like nonlensed events. Its complementary cumulative distribution function is plotted in Fig. 4. 8% of nonlensed events favor lensing, $\log_{10}\mathcal{B}_{\mathrm{lens}}>0$, with a tail that extends to large Bayes factors. We find that GW231123 has a $4\sigma$ false-alarm probability. This is to be contrasted with a set of 1,000  GW231123-like lensed simulations. There, 58% of events have larger Bayes factor than GW231123. Despite the clear difficulties to achieve high detection statistics in events like GW231123, we project that $\sim$40% of our lensed simulations would favor lensing with a significance larger than $5\sigma$.
@@
Figure 4: Complementary cumulative distribution function of lensing Bayes factors ($\mathcal{B}_{\mathrm{lens}}$) for simulated GW231123-like events. The $\log_{10}\mathcal{B}_{\mathrm{lens}}$ of GW231123, 4.0, is indicated by the vertical dashed line. The shaded region corresponds to the Poisson error $1/\sqrt{N}$. Small vertical lines indicate the Bayes factors of the examples displayed in Fig. 2.
@@
Taking advantage of recent advances in GW parameter estimation using deep learning with DINGO [40, 41, 42], we develop DINGO-lensing to accelerate inferences for lensing by orders of magnitude. We exploit DINGO-lensing to thoroughly reanalyze GW231123, the highest ranked lensed candidate to date. However, we find that the false alarm probability of the apparent support for lensing is around $4\sigma$, and could be explained by the self-similarity of the signal. Other potential sources of false evidence of lensing include waveform systematics and non-stationary noise features [31]. That said, these would only increase the false alarm probability for the support. Our method thus gives us the upper bound on the statistical significance in this scenario.
@@
Although we have focused our analyses on searching for lensed signals, the methods and strategies presented here are equally applicable to the identification of waveform modifications from other scenarios, e.g. modified gravity or environmental effects. Our results demonstrate that deep-learning accelerated analysis will play a key role in the discovery of new phenomena in GW astronomy.

### v2

@@ Abstract
Gravitational waves (GWs) are unique messengers as they travel through the Universe without alteration except for gravitational lensing. Their long wavelengths make them susceptible to diffraction by cosmic structures, providing an unprecedented opportunity to map dark matter substructures. Identifying lensed events requires the analysis of thousands to millions of simulated events to reach high statistical significances. This is computationally prohibitive with standard GW parameter estimation methods. We exploit DINGO-lensing, a deep-learning algorithm that accelerates the inference from CPU days to minutes to thoroughly reanalyze GW231123, the most promising lensing candidate to date. By performing more than 200,000 simulations with 3 different waveform models, we find that its statistical significance is below $4\sigma$ and the event cannot be claimed as lensed. We observe that 8% of GW231123-like nonlensed simulations favor lensing, which could be explained by the self-similarity of short-duration signals. Still, 58% of GW231123-like lensed simulations have larger support for lensing, showing that higher detection statistics are possible. We show that analyzing simulations with different waveform models only lowers the significance, highlighting the relevance of waveform systematics. Although GW231123 exposes the challenges of claiming the first GW lensing detection, our deep-learning methods have demonstrated to be powerful enough to enable the upcoming discovery of lensed GWs.
@@
GW parameter estimation has recently experienced a revolution, thanks to deep-learning techniques. In particular, neural posterior estimation [40, 41, 42] is able to drastically reduce the inference time from order of days down to minutes by avoiding the expensive likelihood evaluations needed by conventional methods [43]. This capability enables the assessment of statistical significance for subtle astrophysical effects, such as orbital eccentricity [44], which was previously not feasible due to the prohibitive computational cost of traditional sampling algorithms. In this Letter, we demonstrate that deep-learning techniques are essential for the discovery of lensed GWs. Building on top of state-of-the-art algorithm, DINGO [40, 42, 45, 46, 47], we developed DINGO-lensing [48], which we exploit to thoroughly reanalyze GW231123 and assess its statistical significance for the first time  [^2] Other implementations of lensed parameter estimation with deep-learning have been proposed in parallel [81, 82].. Our results and methods lay out the detections strategies of lensed GWs for upcoming observing runs.
Gravitational waveform distortions from lensing.—Starting from an emitted GW signal $h(t)$, the observed lensed waveform $h_{\rm L}(t)$ is obtained by propagating $h(t)$ through the curved spacetime defined by the lens. Assuming linear perturbations, weak-field gravity, and a lens much smaller than the source–observer distance, this propagation is conveniently solved in Fourier space. Lensing acts like a transfer function between the frequency-domain emitted $\tilde{h}(f)$ and observed signals $\tilde{h}_{\rm L}(f)$, $F(f)=\tilde{h}_{\rm L}(f)/\tilde{h}(f)$, where the amplification function $F(f)$ is defined by the Kirchhoff diffraction integral [24]:
@@
which integrates the time delay $t_{d}$ accumulated by the lensed signal over all paths from source to observer and accounts for their possible interference and diffraction. Here $\vec{\theta}_{\mathrm{S}}$ is the source position, $\vec{\theta}$ the angular position on the lens plane, and $\tau_{D}=\frac{D_{\rm L}D_{\rm S}}{cD_{\rm LS}}$ is a reference time related to the angular diameter distances to the lens $D_{\rm L}$, the source $D_{\rm S}$, and between the lens and the observer $D_{\rm LS}$, respectively. The frequency dependence of $F(f)$ produces the characteristic waveform distortions in $h_{\rm L}(t)$.
The nature of these distortions depends on the lens profile and the source-lens configuration. Recent studies have developed efficient numerical methods to solve the diffraction integral for different lens classes [50, 51, 52]. Observationally, the strongest distortions arise when repeated chirps from a strongly lensed GW source interfere [26, 53], while in the deep wave-optics regime diffraction effects are mild, with $\lim_{f\to 0}F(f)=1$ [24]. This interference regime can be modeled accurately using the stationary phase approximation (SPA) [53], yielding
@@
where $\Delta t\$and $\mu_{\rm rel}\$are the time delay and relative magnification between two repeated signals, respectively. Focusing on the most likely lensing configurations, we assume the second chirp has a lower magnification and opposite parity [54], i.e. $\Delta t\$>0>0, $\mu_{\rm rel}\$<1<1, and a $\pi/2$ phase difference  [^3] Our parametrization differs from the phenomenological model in Ref. [83], where the number of repeated chirps and their phase differences are treated as free parameters.. The overall magnification is omitted, as it is degenerate with luminosity distance. This parametrization encompasses point and fold singularities, as well as many near-cusp configurations [26, 53, 56]. Folds and cusps are the most common caustics in generic lenses [54], while point singularities appear in axisymmetric lenses such as point-mass lenses [57] or Navarro-Frenk-White profiles [58] that can be associated with compact lenses and dark matter subhalos, respectively. Moreover, this simple parametrization has been shown to efficiently recover simulated signals lensed by point-mass lenses [48]. Negative-parity single images, which may cause small additional distortions [59, 60], are excluded here and left for future work.
Under these assumptions, a GW is identified as lensed when the inference yields $\mu_{\rm rel}\$>0>0 and $\Delta t\$>0>0. More complex (and rare) cases beyond Eq. (2), such as overlapping images of the same type with a phase different from $\pi/2$ or $\mu_{\rm rel}\$>1>1 due to substructure [13, 14], will be explored in follow-up work.
@@
We train the lensed networks with $\Delta t\subset(0,0.1)$ sec, and $\mu_{\text{rel}}\subset(0,1)$. The nonlensed parameters follow standard choices for the extrinsic parameters and spins (with magnitudes up to 0.99) [48]. In total there are 17 parameters, which is the largest dimensionality achieved with DINGO so far. Lensed waveforms in DINGO-lensing and bilby are implemented through modwaveforms  [^12] https://github.com/ezquiaga/modwaveforms. We validate the networks as described in [48], testing the recovery of the injected values, comparing the inferences for representative examples with bilby [43], and checking the sampling efficiencies when performing importance sampling [47]. DINGO-lensing  drastically reduces the computing time compared to conventional methods, sampling with neural posterior estimations in seconds and computing evidences with importance sampling in minutes. In comparison, point-mass lens analyses of real events with Gravelamps [50], a code used by LVK, take on average $\sim 50$ CPU days.
Figure 1: Source characterization of GW231123 assuming it is lensed (with DINGO-lensing, in blue) and not lensed (with DINGO, in green), respectively. From the 17 parameters in the inference, we present the chirp mass $\mathcal{M}_{c}$, spin magnitudes $a_{1,2}$, lensing relative magnification $\mu_{\text{rel}}$ and time delay $\Delta t$. The posterior distributions using DINGO-lensing and DINGO are compatible with those obtained using bilby. The credible levels quoted here are from the DINGO-lensing analysis.
@@
Figure 2: Inference for the lensing relative magnification $\mu_{\text{rel}}$ and time delay $\Delta t$. We compare GW231123 results (filled contours) with nonlensed simulations for similarly heavy binaries (unfilled contours).
For the first time, we train neural networks to learn (non)lensed signals using a numerical relativity surrogate waveform model, NRSur7dq4 [69], with a frequency range of $20,512]~\rm{Hz}$, and an analysis duration of $8~\rm{sec}$  [^13] We set the waveform starting frequency at 0 Hz. This means that the model will generate waveforms as far back in time from the merger as allowed.. NRSur7dq4 is the waveform model with the lowest mismatch compared to some GW231123-like numerical relativity simulations [30]. We train the networks for chirp masses in the range of $30,180]~M_{\odot}$, mass ratios in $[1/6,1]$ and luminosity distances in $0.5,10]~\rm{Gpc}$. The prior distribution of the remaining source parameters follow Table 1 in [48], with extended constraints in the component masses. We use the average noise power spectral density of the LIGO detectors around the time of the event [71], and stop the training at epoch 300 when the loss function plateaus.
@@
Figure 3: Whitened best-fit lensed ($h_{\rm L}$, solid), unlensed ($h_{\rm U}$, dash), and nonlensed ($h$, dotted) time-domain waveforms ($\sigma$) for GW231123. The inferred time delay between the two chirps $\Delta t$ matches with their instantaneous period near the reference time $t_{\rm ref}$.
We observe that in many other instances, the recovered $\Delta t$ is multimodal, and the peaks correlate with the period of the waveform. This is due to the self-similarity of short waveforms such as GW231123, for which only a few cycles are observable. We illustrate this using GW231123 in Fig. 3, where we show the inferred whitened best-fit lensed, unlensed (i.e., undoing the lensing amplification from the lensed waveform), and nonlensed (i.e., assuming that there is no lensing) waveforms, respectively  [^15] The whitening uses the estimated noise power spectral density for GW231123 from the LIGO Hanford detector [30].. Our inferred time delay between the two chirps $\Delta t\approx 22$ ms matches with the instantaneous period of the waveforms near the merger at around the reference time $t_{\rm ref}$.
Furthermore, we compare GW231123 Bayes factor with a set of more than 70,000  GW231123-like nonlensed events. Its complementary cumulative distribution function is plotted in Fig. 4  [^16] We have validated representative examples of Fig. 4 computing $\log_{10}\mathcal{B}_{\mathrm{lens}}$ with standard sampling methods (bilby), finding a good agreement (within $\lesssim 0.5$).. 8% of nonlensed events favor lensing, $\log_{10}\mathcal{B}_{\mathrm{lens}}>0$, with a tail that extends to large Bayes factors. We find that GW231123 has a $4\sigma$ false-alarm probability. This is to be contrasted with a set of 1,000  GW231123-like lensed simulations. There, 58% of events have larger Bayes factor than GW231123. Despite the clear difficulties to achieve high detection statistics in events like GW231123, we project that $\sim$40% of our lensed simulations would favor lensing with a significance larger than $5\sigma$.
[…]
@@
Figure 4: Complementary cumulative distribution function $P(x>X)$ of the logarithmic lensing Bayes factors ($x=\log_{10}\mathcal{B}_{\mathrm{lens}}$) for GW231123-like simulations. The $\log_{10}\mathcal{B}_{\mathrm{lens}}$ of GW231123, 4.0, is indicated by the vertical dashed line. The shaded region corresponds to the Poisson error of NN simulations, $1/\sqrt{N}$. Small vertical lines indicate the Bayes factors of the examples displayed in Fig. 2 and the bars on top highlight the false alarm probability levels of the nonlensed simulations in units of standard deviation ($\sigma$). With our neural networks trained on NRSur7dq4, we analyze nonlensed simulations generated with NRSur7dq4 (NRSur), SEOBNRv5PHM (SEOB), and IMRPhenomXPHM (XPHM) waveform approximants..
@@
Taking advantage of recent advances in GW parameter estimation using deep learning with DINGO [40, 41, 42], we exploit DINGO-lensing [48] to thoroughly reanalyze GW231123, the highest ranked lensed candidate to date. We find that the event cannot be claimed as lensed: the false alarm probability of the apparent support for lensing is bounded below $4\sigma$, and could be explained by the self-similarity of the signal. We show that simulating signals with different waveform approximants only lowers the significance, highlighting the relevance of waveform systematics.
@@
Although we have focused our analyses on searching for lensed signals, the methods and strategies presented here are equally applicable to the identification of waveform modifications from other scenarios, e.g. modified gravity environmental effects or overlapping signals. Our results demonstrate that deep-learning accelerated analysis will play a key role in the discovery of new phenomena in GW astronomy.